OpenStudy (anonymous):

For what value of k will x+(k/x) have a relative maximum at x=-2?

OpenStudy (anonymous):

f'(x) = d/dx x + kx^(-1) f'(x) = 1 - kx^(-2) f'(x) = 1 - k/x^2 Then, we can set this equal to zero to get: 1 - k/x^2 = 0 x^2-k = 0 Since x = -2: (-2)^2 - k = 0 4 - k = 0 k = 4